Air is escaping from a spherical balloon at the rate of 2π‘π‘š3 per minute. How fast is the radius shrinking when the volume is 36πœ‹ π‘π‘š3 ?


Find the rate of change of the area 𝐴, of the circle with respect to its circumference C, 𝑖. 𝑒 𝑑𝐴 𝑑�



1
Expert's answer
2021-04-14T12:39:26-0400

for a spherical balloon;

volume, "V= \frac {4}{3}\pi r^3 ...(i)"

area, "A= 4\pi r^2 ...(ii)"

circumference,"C=2\pi r ...(iii)"

a) given "\frac{dv}{dt}=-2 cm^3/min" the negative sign to denote reduction in volume with time and "V= 36\pi cm^3" we have;

"\frac {dv}{dt}=4\pi r^3 \frac {dr}{dt} from (i)"

"\therefore \frac{dr}{dt}= \frac{1}{4\pi r^2} \frac {dv}{dt}...(iv)"

when "V=36\pi\implies \frac {4}{3}\pi r^3 =36\pi"

"r^3= \frac {3*36}{4}=27"

"\implies r=3cm"

"\therefore \frac{dr}{dt}=\frac {1}{4\pi (3)^2}cm/min =\frac{-1}{18\pi}cm/min"

the radius is shrinking at the rate of "\frac{1}{18\pi}cm/min"

b) from (iii) "r= \frac{c}{2\pi}...(v)"

now (v) in (ii) gives;

"A= 4\pi (\frac {c}{2\pi})^2 = \frac {4\pi c^2}{4\pi ^2}= \frac 1\pi c^2 ...(vi)"


"\therefore \frac {dA}{dC}= \frac {2C}{\pi}"


when "V=36\pi cm^3, r= 3cm" Β [from part (a)]

and thus "C=6\pi cm"

"\implies \frac {dA}{dC}= \frac {2(6\pi)} {\pi} (when c=6\pi)"

=12